OpenStudy (anonymous):

Choose the equation of the line passing through the point (-3, 1) and parallel to y = -x - 4. y = x + 2 y = -x - 2 y = -x + 4 y = x - 4

OpenStudy (anonymous):

use the (y-y1)=m(x-x1)

OpenStudy (anonymous):

y-1=-[(-3)-x]

OpenStudy (anonymous):

y-1=-(x+3) y-1=-x-3 add 1 y=-x-2

OpenStudy (anonymous):

Okay??

OpenStudy (anonymous):

does that make sense? you answer is y=-x-2